OpenStudy (anonymous):

Given that f(x) = x2 + 3x + 6 and g(x) = the quantity of x plus two, over three, solve for f(g(x)) when x = 1.

OpenStudy (anonymous):

have you tried it yet?

OpenStudy (anonymous):

No, I dont even know where to start.

OpenStudy (anonymous):

what is equation of g(x) ?

OpenStudy (anonymous):

x-3/2

OpenStudy (anonymous):

(x plus 2) over (3)\[\frac{x+2}{3}\]

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

we want to evaluate f(g(x)) when x=1

OpenStudy (anonymous):

plug in g(x) for every x in f(x) and solve

OpenStudy (anonymous):

so let's find the equation of f(g(x))

OpenStudy (anonymous):

plugging g(x) in f(x) \[f(g(x))=(g(x))^2+3g(x)+6\]and we have the formula for g(x) so\[f(g(x))=\left( \frac{x+2}{3} \right)^2+3\frac{x+2}{3}+6\]

OpenStudy (anonymous):

am i clear? makes sense?

OpenStudy (anonymous):

yeah i think im getting it

OpenStudy (anonymous):

very good, can you evaluate f(g(1)) ?

OpenStudy (anonymous):

oh jeez...haha okay. so is the answer 11

OpenStudy (anonymous):

10 or 11? :-)

OpenStudy (anonymous):

check it once more :-)

OpenStudy (anonymous):

10

OpenStudy (anonymous):

good

OpenStudy (anonymous):

oh i got it!! thank you so much!!!

OpenStudy (anonymous):

Welcome :-)